s. Here are simplified financial statements for Phone Corporation in a recent year: INCOME STATEMENT Figures in miltions ol dolan) Cost of goods sol Other expenses 4060 049 tarmings befone intenet and taxes0T 2564 nterest expense before tax aeat 359 1223 ALANCE SHEET Receivable Oter long-term ansels Total asset Labilities and shareholders eaut 77340 Payables Calculate the following financial ratios for Phone Corporation: a Return on assets (Use after-tax operating income and average assets) Inventory turnover (Use beginning inventory) Long-term debt ratio (Use end of year values) Quick ratio (Use end of year values) e Average collection period (Use beginning receivables) (two points each)
Answer #1

a. Return on assets = After-tax operating income/Average assets = $1668/$27608.50 = 6.04%

After-tax operating income = $2566 - (35% x $2566) = $2566 - $898 = $1668

Average assets = ($27714 + $27503)/2 = $55217/2 = $27608.50

b. Inventory turnover = Cost of goods sold/Beginning inventory = $4060/$238 = 17.06 times

c. Long-term debt ratio = Long-term debt/Total assets = $13196/$27714 = 47.61%

Long-term debt = $7018 + $6178 = $13196

d. Quick ratio = Quick assets/Current liabilities = $2471/$4794 = 0.52 to 1

Quick assets = Cash and marketable securities + Receivables = $89 + $2382 = $2471

e. Average collection period = 365 days/Receivables turnover

Receivables turnover = Net sales/Beginning receivables = $13193/$

Note: Kindly provide the beginning receivables to complete e. since the same is not clear in the image provided.

Note: All answers have been rounded off to 2 decimal places in the absence of specific instructions regarding the same. Kindly round off as required.